### v1

@@ Abstract
Quantum circuits must be error-resilient, yet LLMs like Granite-20B-Code and StarCoder often output flawed Qiskit code. We fine-tuned a 32 B model with two RL methods, Group Relative Policy Optimization (GRPO) and Odds-Ratio Preference Optimization (ORPO), using a richly annotated synthetic dataset. On the Qiskit HumanEval benchmark, ORPO reaches 56.29% Pass@1 ($\approx+10$ pp over Granite-8B-QK) and GRPO hits 49%, both beating all general-purpose baselines; on the original HumanEval they score 65.90% and 63.00%. GRPO excels on basic tasks (42/54), ORPO on intermediate ones (41/68), and neither solves the five advanced tasks, highlighting clear gains yet room for progress in AI-assisted quantum programming.
@@ 1 Introduction
To bridge this gap, our work proposes a Qiskit-based quantum computing coding assistant, an AI-driven tool designed to help developers write and refine quantum programs using the Qiskit SDK. Our assistant focuses specifically on Qiskit, IBM’s widely used quantum SDK, and supports tasks such as circuit construction, optimization, and code debugging. We envision a system that can understand high-level intentions (e.g., “prepare a Bell state” or “optimize this circuit section”) and provide context-sensitive suggestions or code snippets, much like modern code assistants do for classical languages. Using a large language model trained on quantum programming data, the assistant can generate Qiskit code, recommend quantum algorithmic patterns, and catch common mistakes, all within the developer’s workflow. The goal is to lower the barrier to entry for quantum programming and to accelerate the development process for experts and beginners alike.
@@ 2.1 Quantum Programming Environments and Tools
In addition to Qiskit, practical resources such as Silva (2018) provide essential hands-on strategies for programming quantum rigs using Python, the quantum assembly language, and cloud-based platforms, including IBM QExperience [8]. This resource highlights the inherent challenges in quantum programming and emphasizes the need for developer-friendly methodologies and tools.
@@ 2.2 AI-Assisted Quantum Code Generation
By positioning our assistant within existing quantum programming environments and AI coding tools, we aim to advance the support available to quantum computing developers. Ultimately, this work contributes to the broader goal of quantum software engineering: to bring the productivity and reliability benefits of modern software development to the field of quantum computing [5], thereby accelerating innovation and adoption.
@@ 3.1 Generation of Training Data
To enable robust supervised and reinforcement learning, we construct a high-quality dataset comprising 522 Qiskit programming tasks. The data set was created through a multistage pipeline that includes code retrieval, function extraction, annotation, validation, deduplication, and formatting. The entire process is automated and designed to ensure consistency, reproducibility, and broad coverage of real-world quantum programming challenges.
@@
To improve training diversity and reduce redundancy, we apply structural and semantic deduplication techniques. Near-duplicate solutions or trivial variants are filtered out using syntactic similarity and abstract syntax tree (AST) comparisons, ensuring a more diverse set of training signals.
@@ 3.2 Reinforcement Learning with Preferences
$\text{KL}(P\|Q)=\sum_{i}P(i)\log\left(\frac{P(i)}{Q(i)}\right)$
@@
$\mu=\frac{1}{G}\sum_{i=1}^{G}r(y_{i}),\quad\sigma=\sqrt{\frac{1}{G}\sum_{i=1}^{G}(r(y_{i})-\mu)^{2}},\quad A(y_{i})=\frac{r(y_{i})-\mu}{\sigma}.$
@@
The clip function $\operatorname{clip}(v,L,U)=\max(L,\min(v,U))$ bounds $r_{\theta}$, and $\epsilon$ (e.g., 0.10.10.10.10.20.20.20.2) sets the range $[1-\epsilon,1+\epsilon]$. This process guides the model toward producing more executable and resource-efficient quantum circuits by emphasizing outputs that outperform others in the same generation group.
@@ 4.1 Evaluation Setup
Although the QHE paper reports 101 evaluation tasks, the publicly released dataset contains 151 entries, many of which are duplicates or auxiliary examples. We filter the data set to retain 101 uniquely identifiable tasks and group them by difficulty based on metadata from the original authors.
Since the original evaluation script was not publicly released and the HumanEval [15] framework was incompatible with QHE tasks, we implemented a custom benchmarking script customized for the QHE setting. This script automatically executes model completions against the associated unit tests and logs pass/fail outcomes, enabling consistent and scalable evaluation across all models. Our results are therefore based on a fully automated, reproducible evaluation pipeline that faithfully adheres to the QHE benchmark structure.
We compared our GRPO and ORPO models against several strong baseline models:
@@
Table 2 summarizes the hyperparameter settings used during fine-tuning for both approaches that we ran on the A100 GPU with 80 GB for RAM.
@@ 4.2 Results
Table 3: Pass@1 on HumanEval (HE) and Qiskit HumanEval (QHE) with Greedy Decoding
@@
Table 4: QHE Pass Counts by Difficulty Level (54 Basic, 45 Intermediate, 2 Advanced)
@@
To better understand the behavior of the model across the complexity of tasks, Table 4 presents pass counts grouped by difficulty level. GRPO leads in basic-level tasks with 42/54 passes and shows robust performance on intermediate tasks (32/68). ORPO achieves the highest pass count on intermediate tasks (41/68) and performs slightly better overall in total completions. Neither model succeeds on the five advanced tasks, consistent with all other baselines.
@@
In particular, even official benchmark completions passed only 69/78 Basic, 63/68 Intermediate, and 2/5 Advanced tasks when evaluated with our benchmarking script, suggesting that changes in the Qiskit version may have impacted the stability of the evaluation. Nevertheless, both models perform on par with or exceed these reference rates, further validating their practical utility.
Overall, the results highlight the strength of preference-driven optimization in quantum programming and emphasize the importance of evaluating across difficulty levels to capture nuanced model capabilities.  [^1] The ’out-of’ counts for our models reflect the number of unique and valid tasks from the publicly released Qiskit HumanEval dataset we were able to evaluate using our custom benchmarking script, which differed from the task counts reported in the original QHE paper due to dataset inconsistencies.
@@ 4.2.1 Training Dynamics
Figure 1: Training dynamics of our fine-tuned models. Left: Reward fluctuations during GRPO optimization. Right: Loss convergence during ORPO fine-tuning.
@@ 4.3 Discussion
Our updated evaluation highlights the effectiveness of reinforcement learning with preferences, through GRPO and ORPO, for quantum code generation. Both models substantially outperform general-purpose LLMs on the Qiskit HumanEval (QHE) benchmark under greedy decoding. ORPO achieves the highest QHE pass@1 score at 56.29%, surpassing even the domain-specific Granite-8B-Code-QK baseline (46.53%), while GRPO also delivers a strong 49.00%. These improvements were achieved without explicit supervised instruction tuning or access to the original QHE fine-tuning scripts, emphasizing the strength of our preference optimization pipeline.
At the task difficulty level, GRPO and ORPO demonstrate complementary advantages. GRPO achieves the highest pass count in basic tasks (42/54), suggesting that its group-based optimization effectively promotes structural correctness and diversity in simpler circuits. ORPO, on the contrary, leads on intermediate tasks (41/68), indicating that learning from ranked preferences better captures subtleties such as readability, robustness, and correctness in more complex quantum programs. These trends suggest that combining GRPO and ORPO may yield even stronger performance through hybrid reward strategies.
Interestingly, even the benchmark reference implementation, executed via the latest version of Qiskit, only achieves 69/78 on Basic, 63/68 on Intermediate, and 2/5 on Advanced tasks. This highlights potential fragility and version sensitivity in quantum execution environments. It also raises a critical point: our models were tested under realistic run-time conditions and still matched or exceeded these reference pass counts, reinforcing their practical applicability.
Compared to previous work by Vishwakarma et al. [3], our results are competitive despite several evaluation challenges. Although their paper reports 101 tasks, the public release contained 151 files, many of which were duplicates or auxiliary entries. Furthermore, their evaluation script was not released and the adaptation of the HumanEval framework failed to validate even official completions. To address this, we wrote our benchmark evaluation script and validated completions using the unit tests provided, ensuring consistent and reproducible results.
@@ 5 Conclusion and Future Work
However, this work also presents important challenges. We encountered inconsistencies in benchmark releases, missing evaluation scripts, and had to manually run and validate test cases, which affected reproducibility. Furthermore, none of the evaluated models, including ours, succeeded in the Advanced-level tasks, pointing to the need for better instruction tuning, longer-horizon reasoning, and deeper integration with quantum hardware constraints. Although our results demonstrate the potential for preference-based optimization for quantum code generation, there are several limitations to be acknowledged. First, the training data remains relatively small compared to classical code datasets, which may restrict generalization to novel quantum tasks. Second, due to the absence of an official evaluation script and inconsistencies in the published benchmark, we relied on manual validation for scoring, introducing potential subjectivity, and making direct comparisons to other models less precise. Finally, the current models have not yet been tested on real hardware, and their robustness under execution noise remains unexplored.

### v2

@@ Abstract
Quantum circuits must be error-resilient, yet LLMs like Granite-20B-Code and StarCoder often output flawed Qiskit code. We fine-tuned the Qwen2.5-Coder-32B model with two RL methods, Group Relative Policy Optimization (GRPO) and Odds-Ratio Preference Optimization (ORPO), using a richly annotated synthetic dataset. On the Qiskit HumanEval benchmark, ORPO reaches 56.29% Pass@1 ($\approx+10$ pp over Granite-8B-QK) and GRPO hits 49%, both beating all general-purpose baselines; on the original HumanEval they score 65.90% and 63.00%. GRPO performs well on basic tasks (44/78) and excels on intermediate ones (41/68), but neither GRPO nor ORPO solves any of the five advanced tasks, highlighting clear gains yet room for progress in AI-assisted quantum programming.
@@ 1 Introduction
To bridge this gap, our work proposes a Qiskit-based quantum computing coding assistant, an AI-driven tool designed to help developers write and refine quantum programs using the Qiskit SDK. QSpark [^1] To facilitate reproducibility, we release our full implementation at https://github.com/TMUDeV/QSPARK. focuses specifically on Qiskit, IBM’s widely used quantum SDK, and supports tasks such as circuit construction, optimization, and code debugging. We envision a system that can understand high-level intentions (e.g., “prepare a Bell state” or “optimize this circuit section”) and provide context-sensitive suggestions or code snippets, much like modern code assistants do for classical languages. Using a large language model trained on quantum programming data, the assistant can generate Qiskit code, recommend quantum algorithmic patterns, and catch common mistakes, all within the developer’s workflow. The goal is to lower the barrier to entry for quantum programming and to accelerate the development process for experts and beginners alike.
@@ 2.1 Quantum Programming Environments and Tools
In addition to Qiskit, practical resources such as the book by Silva [8] provide essential hands-on strategies for programming quantum rigs using Python, the quantum assembly language, and cloud-based platforms, including IBM QExperience. This resource highlights the inherent challenges in quantum programming and emphasizes the need for developer-friendly methodologies and tools.
@@ 2.2 AI-Assisted Quantum Code Generation
By positioning QSpark within existing quantum programming environments and AI coding tools, we aim to advance the support available to quantum computing developers. Ultimately, this work contributes to the broader goal of quantum software engineering: to bring the productivity and reliability benefits of modern software development to the field of quantum computing [5], thereby accelerating innovation and adoption.
@@ 3.1 Generation of Training Data
To enable robust supervised and reinforcement learning, we construct a high-quality dataset, for training purposes, comprising 522 Qiskit programming tasks. The data set was created through a multistage pipeline that includes code retrieval, function extraction, annotation, validation, deduplication, and formatting. The entire process is automated and designed to ensure consistency, reproducibility, and broad coverage of real-world quantum programming challenges.
@@
Table 1: Difficulty scale with representative Qiskit programming tasks.
Level Criteria and Example Task
Basic Simple circuits with few gates and minimal measurement. Example: prepare a single qubit in superposition using an H-gate.
Intermediate Circuits with measurement, moderate depth, or algorithmic structure. Example: implement a 4-qubit Quantum Fourier Transform (QFT).
Difficult Complex circuits involving entanglement, variational methods, or hybrid workflows. Example: build a Variational Quantum Eigensolver (VQE) ansatz circuit and connect it to a classical optimizer.
@@
To improve training diversity and reduce redundancy, we apply structural and semantic deduplication techniques. Near-duplicate solutions or trivial variants are filtered out using syntactic similarity and abstract syntax tree (AST) comparisons, ensuring a more diverse set of training signals. To illustrate the filtering process, approximately 10,819 raw Qiskit-related functions were initially collected, of which fewer than 5% passed all validation steps. The majority of rejections were due to incomplete docstrings, missing test coverage, or trivial circuits (e.g., functions that only returned an empty register). The final curated set of 522 tasks was therefore deliberately biased toward code that was both executable and semantically meaningful. Difficulty scores were computed automatically using heuristics: basic tasks had depth $\leq 3$ and no entanglement; intermediate tasks required either measurements or circuit depth >3>3; and difficult tasks contained multi-qubit entanglement or hybrid classical-quantum structures. Unlike the QHE benchmark, which evaluates generalization, this dataset was designed for training, and thus prioritizes diversity and coverage across circuit patterns.
@@ 3.2 Reinforcement Learning with Preferences
$\text{KL}(P\|Q)=\sum_{i}P(i)\log\left(\frac{P(i)}{Q(i)}\right)$ (2)
@@
ORPO Reward Construction.
Odds-Ratio Preference Optimization (ORPO) aligns the model with human-like coding preferences, focusing on readability and maintainability. For each prompt, we construct a pairwise comparison between a chosen output $\hat{y}$ and a rejected output $y$. The chosen output is correct, executable, and stylistically aligned with Qiskit best practices. while the rejected output is either synthetically perturbed or sampled from lower-quality generations.
@@
$\mu=\frac{1}{G}\sum_{i=1}^{G}r(y_{i}),\quad\sigma=\sqrt{\frac{1}{G}\sum_{i=1}^{G}(r(y_{i})-\mu)^{2}},\quad A(y_{i})=\frac{r(y_{i})-\mu}{\sigma}.$ (3)
@@
The clip function $\operatorname{clip}(v,L,U)=\max(L,\min(v,U))$ bounds $r_{\theta}$, and $\epsilon$ (e.g., 0.10.10.20.2) sets the range $1-\epsilon,1+\epsilon]$. This process guides the model toward producing more executable and resource-efficient quantum circuits by emphasizing outputs that outperform others in the same generation group.
GRPO Reward Construction.
For each prompt, we generate a group of candidate completions. Each candidate is executed with Qiskit v2.0.0 and Qiskit Aer v0.17.1 simulators. The execution is scored using three criteria:
  1. 1.
    Unit test pass rate ($r_{1}$): fraction of test cases passed (primary correctness signal).
  2. 2.
    Circuit depth penalty ($r_{2}$): normalized inverse of circuit depth to reward more efficient solutions.
  3. 3.
    Qubit count penalty ($r_{3}$): normalized inverse of the number of qubits used, discouraging wasteful allocations.
These are combined into a single scalar reward:
$r(y)=\alpha\cdot r_{1}+\beta\cdot r_{2}+\gamma\cdot r_{3},$ (5)
with weights $\alpha=0.7$, $\beta=0.2$, and $\gamma=0.1$ chosen empirically to emphasize correctness while still encouraging efficiency.
Within each group, rewards are normalized using the group mean and variance:
$A(y_{i})=\frac{r(y_{i})-\mu}{\sigma},\quad\mu=\frac{1}{G}\sum_{i=1}^{G}r(y_{i}),\quad\sigma=\sqrt{\frac{1}{G}\sum_{i=1}^{G}(r(y_{i})-\mu)^{2}},$ (6)
where $G$ is the number of candidates. This normalization ensures that the rewards are relative: a candidate only receives a high advantage if it is better than its peers, even if all solutions are weak. The policy update (Eq. 4) then uses this normalized advantage to push the model towards consistently producing correct and efficient circuits.
In practice, this setup allows GRPO to prefer structurally sound and resource-efficient quantum circuits for simple tasks, while avoiding overfitting to a single absolute scoring heuristic.
@@ 4.1 Evaluation Setup
Since the original evaluation script was not publicly released and the HumanEval [15] framework was incompatible with QHE tasks, we implemented a custom benchmarking script customized for the QHE setting. Although the QHE paper reports 101 tasks, the publicly released dataset contains 151 entries. This resulted in 78 Basic, 68 Intermediate, and 5 Advanced tasks. This script automatically executes model completions against the associated unit tests and logs pass/fail outcomes, enabling consistent and scalable evaluation across all models. Our results are therefore based on a fully automated, reproducible evaluation pipeline that faithfully adheres to the QHE benchmark structure.
We compared our GRPO and ORPO models with the following strong baseline models.
@@
Table 3 summarizes the hyperparameter settings used during fine-tuning for both approaches that we ran on the A100 GPU with 80 GB for VRAM.
@@
Batch Size 64 32
Sequence Length 2048 tokens 2048 tokens
@@ 4.2 Results
Table 4: Pass@1 on HumanEval (HE) and Qiskit HumanEval (QHE) with Greedy Decoding. The results for models other than ours are taken from [3].
@@
Table 5: QHE Pass Counts by Difficulty Level (78 Basic, 68 Intermediate, 5 Advanced). The results for models other than ours are taken from [3].
@@
To better understand the behavior of the model across the complexity of tasks, Table 5 presents pass counts grouped by difficulty level. ORPO ranks third in basic-level tasks with 44/78 (slightly behind Granite-8B-Code-QK at 32/54 and DeepSeek-Coder-33B at 30/54 in terms of completion percentage), achieves the highest pass count on intermediate tasks (41/68), and outperforms others in total completions. GRPO performs worse than ORPO but still surpasses many baseline models. Neither model succeeds on the five advanced tasks, consistent with all other baselines.
@@
Both models perform on par with or exceed these reference rates, further validating their practical utility. Overall, the results highlight the strength of preference-driven optimization in quantum programming and emphasize the importance of evaluating across difficulty levels to capture nuanced model capabilities.
@@ 4.2.1 Training Dynamics
Figure 1: Training dynamics of QSpark fine-tuning. Left: GRPO reward trajectory with normalized advantage values ($y$-axis) across training steps ($x$-axis). Right: ORPO pairwise ranking loss ($y$-axis) across training steps ($x$-axis). Legends and axis labels are explicitly provided to highlight reward and loss behavior.
@@ 4.3 Discussion
Our updated evaluation highlights the effectiveness of reinforcement learning with preferences, through GRPO and ORPO, for quantum code generation. Both models substantially outperform general-purpose LLMs on the Qiskit HumanEval (QHE) benchmark under greedy decoding. ORPO achieves the highest QHE pass@1 score at 56.29%, surpassing even the domain-specific Granite-8B-Code-QK baseline (46.53%), while GRPO also delivers a strong 49.00%. These improvements were achieved without explicit supervised instruction tuning or access to the original QHE fine-tuning scripts, emphasizing the strength of our preference optimization pipeline. At the task difficulty level, ORPO consistently outperforms GRPO.
Interestingly, even the benchmark reference implementation, executed via the latest version of Qiskit, only achieves 69/78 on Basic, 63/68 on Intermediate, and 2/5 on Advanced tasks. This highlights potential fragility and version sensitivity in quantum execution environments. It also raises a critical point: our models were tested under realistic run-time conditions and still matched or exceeded these reference pass counts, reinforcing their practical applicability. Compared to previous work by Vishwakarma et al. [3], our results are competitive despite several evaluation challenges. Although their paper reports 101 tasks, the public release contained 151 files. Furthermore, their evaluation script was not released. To address this, we wrote our benchmark evaluation script and validated completions using the unit tests provided, ensuring consistent and reproducible results.
@@ 5 Conclusion and Future Work
However, this work also presents important challenges. We encountered inconsistencies in benchmark releases, missing evaluation scripts, and had to manually run and validate test cases, which affected reproducibility. Furthermore, none of the evaluated models, including ours, succeeded in the Advanced-level tasks, pointing to the need for better instruction tuning, longer-horizon reasoning, and deeper integration with quantum hardware constraints. Although our results demonstrate the potential for preference-based optimization for quantum code generation, there are several limitations to be acknowledged. First, the training data remains relatively small compared to classical code datasets, which may restrict generalization to novel quantum tasks. Second, due to the absence of an official evaluation script and inconsistencies in the published benchmark, we relied on manual validation for scoring, introducing potential subjectivity, and making direct comparisons to other models less precise.
@@
6 Acknowledgements
This work is partially sponsored by Natural Science and Engineering Research Council of Canada (grants # 2020-04760 and RGPIN-2022-03886).